# SpatialDreamer: Self-supervised Stereo Video Synthesis from Monocular Input

### v1

@@ Abstract
Stereo video synthesis from a monocular input is a demanding task in the fields of spatial computing and virtual reality. The main challenges of this task lie on the insufficiency of high-quality paired stereo videos for training and the difficulty of maintaining the spatio-temporal consistency between frames. Existing methods primarily address these issues by directly applying novel view synthesis (NVS) techniques to video, while facing limitations such as the inability to effectively represent dynamic scenes and the requirement for large amounts of training data. In this paper, we introduce a novel self-supervised stereo video synthesis paradigm via a video diffusion model, termed SpatialDreamer, which meets the challenges head-on. Firstly, to address the stereo video data insufficiency, we propose a Depth based Video Generation module DVG, which employs a forward-backward rendering mechanism to generate paired videos with geometric and temporal priors. Leveraging data generated by DVG, we propose RefinerNet along with a self-supervised synthetic framework designed to facilitate efficient and dedicated training. More importantly, we devise a consistency control module, which consists of a metric of stereo deviation strength and a Temporal Interaction Learning module TIL for geometric and temporal consistency ensurance respectively. We evaluated the proposed method against various benchmark methods, with the results showcasing its superior performance.
@@ 1 Introduction
In this paper, we propose a self-supervised stereo video synthesis paradigm via a video diffusion model, termed SpatialDreamer, which will meet the data insufficiency and spatio-temporal inconsistency at once. Two novel designs make our SpatialDreamer non-trivial. Firstly, we design a Depth based Video data Generation module DVG to handle the problem of data insufficiency. Similar to Wang et al. [64], DVG constructs a training pair of two views by utilizing forward-backward rendering, without the need of data annotation [57, 49, 68]. However, Wang et al. [64] only renders individual images, which will produce flickering results for video data generation [9]. In this paper, we improve the generation of video data for refining stereo occlusion masks by leveraging the inter-frame motion obtained from optical flow [28]. This methodology enables DVG to produce paired video data that maintains both geometric and temporal consistency. Leveraging paired video generated by DVG, we propose RefinerNet, along with a self-supervised video synthesis framework, to enable efficient and targeted training. More importantly, with access to sufficient paired videos, we devise a consistency control module, which consists of a metric of stereo deviation strength and a Temporal Interaction Learning module TIL. The stereo deviation strength aims to enable the generation of stereo videos in diverse real-world scenarios. A stereo-aware loss is further used to supervise the model in learning the magnitude of difference between the paired view features in the latent space. TIL integrates the latent features from long-temporal frames as a global information to augment the temporal coherence of the generated results.
@@
Figure 2: Overview of the proposed method, given left view video, the target view video is rendered, encoded, and concatenated with multi-frame noise, followed by the denoising U-Net architecture (i.e., SVD). The reference view images are fed into RefinerNet, through which the spatial features are extracted. (Notably, the left and right views are the target and reference view respectively for training, and the opposite for inference). The temporal interaction learning module integrates the latent features from long-temporal frames, and the deviation strength is projected as positional embedding and added to the time step embedding. Finally, the variational autoencoder decoder decodes the result into a video clip.
@@ 2.1 Single-image Novel View Synthesis
Early studies address the single-image novel view synthesis task without relying on explicit 3D representation, such as End-to-end Synthesis [67], Infinite-Nature [35] and Animatable 3D Characters [66]. These methods do not utilize explicit 3D representations for entire scenes, which limits the consistency of 3D views in the generated outputs and results in flickering and blurring, particularly in occluded regions. Layer-based methods, such as 3D-photography [54] and SLIDE [26] represent a 3D scene using discrete layers. Yet, the synthesized views may lack accuracy in texture and geometric consistency across different perspectives. SMPI [59], AdaMPI [21] and SinMPI [45] represent the 3D scene using multiple planes, while their rendered results may suffer from depth discretization artifacts and repetitive texture patterns. To reduce the reliance on large labeled datasets, Shi et.al [53], Wang et al. [64] and NVSVDE-Net [4] explore self-supervised learning for creating novel view images from single image. Although they are capable of producing photorealistic results, they struggle to maintain temporal and spatial consistency, particularly in dynamic scenes.
@@ 2.2 Multi-view Novel View Synthesis
Early approaches in this field frequently utilized geometric methods, such as multi-view stereo reconstruction [52, 29] and structure-from-motion techniques [51, 44], to generate novel viewpoints from multiple images captured at different angles. NeRF-based methods were proposed to generate novel views based on fewer input images, including SinNeRF [70], CoPoNeRF [24] and NViST [27] . 3D Gaussian Splatting [31] presented an explicit representation of the scene. Gaussian-based methods, such as pixelSplat [12] and MVSplat [14], can produce high-quality synthesis results from multi-view images. Recently, researchers have begun exploring the application of advanced diffusion models for novel view synthesis, driven by their remarkable success in this area. Notable approaches such as Photoconsistent-NVS [73], MultiDiff [40], GCD [61] and NVS-Solver [72] generate image sequences based on predefined camera trajectories using either single or multiple input images.
While these methods have shown promising results, they often suffer from limitations such as dependence on dense scene geometry, handling complex scene dynamics, and being hard to generalize [31]. Moreover, the challenge of optimizing NeRF/GS with a limited number of input views hampers the ability of multi-view NVS to produce high-quality results.
@@ 3 SpatialDreamer
Given an image sequence and a stereoscopic pose describing the scene depth information, the proposed method generates a stereo pair. In this section, we first provide a brief introduction about SVD [7]. Secondly, we offer detailed information of our proposed video generation module DVG, which can generate sufficient paired video with geometric and temporal consistency. Based on the DVG, we present the framework for self-supervised learning based video synthesis to support efficient and dedicate training. More importantly, to ensure the consistency between frames, we design a consistency control module for geometric and temporal consistency ensurance. Finally, we present the detailed training process.
@@ 3.1 Preliminaries
The challenge in video generation lies in accurately modeling the spatial-temporal dynamics, which involve understanding the spatial relationships within individual frames and the temporal relationships across frames. To address this challenge, Stable Video Diffusion (SVD) [7] introduces temporal convolution and attention layers. SVD follows the latent diffusion model (LDM) to encode video pixels into the latent space, enabling a more efficient denoising process. Given a video $I_{v}$, encoder $\mathcal{E}$ encodes $I_{v}$ into the latent space as $z_{0}=\mathcal{E}(I_{v})$. Gaussian noise $\epsilon$ is then added in the Markov process:
@@
where $\bar{\alpha}_{0},\ldots,\bar{\alpha}_{T}$ are the pre-defined noise schedules for T steps in the Markov process. Notably, SVD is trained by the $v$-prediction formulation with the mean squared error (MSE) between the ground truth and its prediction in the latent space.
@@ 3.2 Depth based Video Generation
To create training pairs of images along with their corresponding occluded masks, the following self-supervised mechanism is employed utilizing real images. Firstly, the depth of the given video sample is estimated by MiDaS [47]. Secondly, the reference view image $x_{1}$ is rendered into a masked one under the target viewpoint $P_{2}$, and then the trained inpainting model [56] is used to fill these occluded regions to obtain the novel viewpoint $x_{2}$ [41]. Finally, $x_{2}$ is rendered back to the original viewpoint $P_{1}$ via a back rendering. After this, we can obtain the masked image $\widetilde{x}_{1}$ with mask $M$ under viewpoint $P_{1}$, as well as the inpainted image $x_{2}$ under viewpoint $P_{2}$. The mask $M$ and masked image $\widetilde{x}_{1}$ are considered to be the conditions, with the image $x_{1}$ serving as the ground truth for the diffusion model. This approach involves encoding the inpainted image $x_{2}$ into latent feature representations and feeding it into RefinerNet.
@@ 3.3 Self-Supervised Stereo Video Synthesis framework
Leveraging the generated paired view videos, we propose a self-supervised framework for stereo video synthesis, as illustrated in Fig 2. This framework commences with the rendering of the target view video that incorporates occlusions, followed by the application of a spatio-temporal method to effectively inpaint the missing regions, ultimately yielding a complete non-occluded target view video.
To learn the distribution of the feature space between the paired view images, we devise RefinerNet, which is modeled after the architecture of the denoising U-Net, excluding the temporal layer. RefinerNet adopts the weight initialization from the original SD2.1 model [48]. As depicted in Figure 2, a spatial attention layer is employed instead of the self-attention layer. Given a feature map $z_{t}\in\mathbb{R}^{t\times h\times w\times c}$ from the denoising U-Net and another feature map $z_{r}\in\mathbb{R}^{t\times h\times w\times c}$ from RefinerNet along with TIL module, they are concatenated along the h dimension, and then self-attention is applied. Subsequently, the first half of the feature map is retrieved as the output. One great advantage of this approach is that it enables RefinerNet to maintain the feature representation capability inherent in the original SD Unet, extracting reference view image information. In addition, the denoising U-Net can adaptively learn features from RefinerNet that are correlated in the same feature space, owing to the fundamentally similar network architectures between the denoising U-Net and RefinerNet. This not only ensures a robust initialization for the feature representation but also enhances the learning ability of the entire network.
@@ 3.4.1 Temporal Interaction Learning module
While our RefinerNet-based approach providing spatial guidance, paired view features are still merged independently, posing challenges for temporal consistency and the generated images may exhibit discrepancies in texture or produce peculiar results, particularly between adjacent frames.
Inspired by previous studies [11, 22], we propose an attention-based temporal interaction learning module that learns the distribution of the feature space from adjacent frames, named TIL. Specifically, given latent codes of $N_{r}$ adjacent images under reference view $z_{i}^{t}$, where $i$ = 1, 2, …, $N_{r}$ and the latent features of the reference view image $z_{r}^{t}$ at the time step $t$, we augment the $z_{r}^{t}$ to $aug_{r}^{t}$ as:
@@
where $\lambda\in[0,1]$ and $\mathrm{Attn}_{i,j}$ defines attention between two latent features $z_{i}$ and $z_{j}$. This module blends the self-attention of $z_{r}^{t}$ with the cross-view attention between $z_{r}^{t}$ and each adjacent view $z_{i}^{t}$. The cross-view attention aligns the feature of all adjacent frames to the reference view while the self-attention helps each reference image retain distinctiveness.
The augmented reference feature $aug_{r}^{t}$ is then fed into spatial attention layer to assist U-Net network learning.
@@ 3.5 Training Strategy
The training process is divided into two stages. In the first training stage, we focus on individual frames from videos. The temporal layer in denoising U-Net is frozen, and TIL module is removed. The RefinerNet model and the denoising U-Net are initialized with pretrained weights from SD2.1 [48] and SVD [7] respectively. The weights of the variational autoencoder encoder and decoder, as well as the contrastive language-image pre-training (CLIP) image encoder, are all kept fixed. The aim of this stage is for the model to learn to synthesize new viewpoint images under the condition of a reference image and a new viewpoint pose. In the second stage, we train the temporal layer and TIL module with video sequences. This enables the model to capture temporal context information efficiently. The input for the model is the 8-frame video clip and $N_{r}$ for TIL is set to 8. During this stage, only the denoising U-Net and TIL module are trained while keeping the weights of the remaining network unchanged.
@@ 4.1 Datasets, Baselines and Metrics
We conduct our experiments on image and video levels with RealEstate10K [76] and self-collected video data. RealEstate10K is a large, widely used dataset with camera poses, corresponding to 10 million frames derived from about 80,000 video clips, gathered from about 10,000 YouTube videos. As there is not open-source stereo video benchmark, we collect 1500 monocular videos with 1400 for training, 100 for testing. For the image synthesis, we compared the proposed method with recent open-source SOTA methods: 3D-photography [54], SynSin [67], AdaMPI [21], SinMPI [45], Wang et al. [64], Photoconsistent-NVS(P-NVS) [73], NVSVDE-Net [4], CoPoNeRF [24] and MVSplat [14]. For a fair comparison, the method of 3D-photography, AdaMPI, SinMPI, and the proposed method all used MiDaS [47] for the depth estimation. All the methods were conducted with the same intrinsic matrices, source camera poses, and target camera poses. Following the method of 3D-photography [54], we measured the accuracy of the generated target views with regard to the ground-truth images using the three metrics of the learned perceptual image patch similarity (LPIPS), peak signal-to-noise ratio (PSNR), and structural similarity index measure (SSIM).
For the video synthesis, we picked several representative approaches, i.e., the method of 3D-photography [54], the open-source code repository webui-depthmap [^1] https://github.com/thygate/stable-diffusion-webui-depthmap-script , P-NVS [73], CoPoNeRF [24], NVS-Solver [72] and AVP [42]. Moreover, we conducted the same depth estimation method [47] for a fair comparison of 3D-photography [54], webui-depthmap and NVS-Solver [72]. We employ the FVD [60] score to measure the perceptual similarity between input videos and outputs and report the flow warping error $E_{warp}$ [34] to assess the temporal consistency of the resulting video sequences.
@@ 4.2 Implementation
We employed the AdamW [37] optimizer with a learning rate of $1\times{10}^{-5}$ for training the model. All the experiments were conducted using two NVIDIA A800 GPUs. In the first training stage, individual video frames were sampled, resized, and center-cropped to a resolution of 1024 × 1024. We trained the model for 100,000 iterations with a batch size of 8. In the second training stage, the videos were first split according to the transitions to ensure that a scene only appeared in one video clip. We then formed each video clip for training with eight frames. Finally, the temporal layer and TIL module was trained for 10,000 steps with a batch size of 2. The stereo-aware loss scaling factor $\lambda$ was set to 0.001 and $\lambda$ in Eq. 3 was set as 0.6.

### v2

@@ Abstract
Stereo video synthesis from monocular input is challenging in spatial computing and virtual reality due to the lack of high-quality stereo video pairs for training and the difficulty of maintaining spatio-temporal consistency between frames. Existing methods primarily address these issues by directly applying novel view synthesis (NVS) techniques to video, while facing limitations such as the inability to effectively represent dynamic scenes and the requirement for extensive training data. In this paper, we introduce a novel self-supervised stereo video synthesis paradigm via a video diffusion model, termed SpatialDreamer, which meets the challenges head-on. Firstly, to address the stereo video data insufficiency, we propose a Depth based Video Generation module DVG, which employs a forward-backward rendering mechanism to generate paired videos with geometric and temporal priors. Leveraging data generated by DVG, we propose RefinerNet along with a self-supervised synthetic framework designed to facilitate efficient and dedicated training. More importantly, we devise a consistency control module, which consists of a metric of stereo deviation strength and a Temporal Interaction Learning module TIL for geometric and temporal consistency ensurance respectively. We evaluated the proposed method against various benchmark methods, with the results showcasing its superior performance. Our project website is at: https://spatialdreamer.github.io.
footnotetext: Corresponding author.
{NoHyper}
@@ 1 Introduction
In this paper, we propose a self-supervised stereo video synthesis paradigm via a video diffusion model, termed SpatialDreamer, which will meet the data insufficiency and spatio-temporal inconsistency at once. Firstly, we design a Depth based Video data Generation module DVG to handle the problem of data insufficiency. Similar to [21, 66] , DVG constructs a training pair of two views using forward-backward rendering without the need for data annotation  [59, 51, 70], while [21, 66] only render individual images, resulting in flickering effects in generated video [9]. In this paper, we improve the generation of video data for refining stereo occlusion masks by leveraging the inter-frame motion obtained from optical flow [29]. This methodology enables DVG to produce paired video data that maintains both geometric and temporal consistency. Leveraging paired video generated by DVG, we propose RefinerNet, along with a self-supervised video synthesis framework, to enable efficient and targeted training. More importantly, with access to sufficient paired videos, we devise a consistency control module, which consists of a metric of stereo deviation strength and a Temporal Interaction Learning module TIL. The stereo deviation strength aims to enable the generation of stereo videos in diverse real-world scenarios, with a stereo-aware loss further supervising the model to learn the magnitude of difference between paired view features in the latent space. TIL integrates the latent features from long-temporal frames as a global information to augment the temporal coherence of the generated results.
@@ 2 Related Work
Figure 2: Overview of the proposed method, given left view video, the target view video is rendered, encoded, and concatenated with multi-frame noise, followed by the denoising U-Net architecture (i.e., SVD). The reference view images are fed into RefinerNet, through which the spatial features are extracted. (Notably, the left and right views are the target and reference view for both training and inference). The temporal interaction learning module integrates the latent features from long-temporal frames, and the deviation strength is projected as positional embedding and added to the time step embedding. Finally, the variational autoencoder decoder decodes the result into a video clip.
@@ 2.1 Single-image Novel View Synthesis
Early studies address the single-image novel view synthesis task without relying on explicit 3D representation, such as End-to-end Synthesis [69], Infinite-Nature [37] and Animatable 3D Characters [68]. These methods lack explicit 3D representations for scenes, leading to inconsistent views and causing flickering and blurring, especially in occluded regions. Layer-based methods, such as 3D-photography [56] and SLIDE [27] represent a 3D scene using discrete layers. Yet, the synthesized views may lack accuracy in texture and geometric consistency across different perspectives. SMPI [61], AdaMPI [21] and SinMPI [47] represent the 3D scene using multiple planes, while their rendered results may suffer from depth discretization artifacts and repetitive texture patterns. To reduce the reliance on large labeled datasets, Shi et.al [55], Wang et al. [66] and NVSVDE-Net [4] explore self-supervised learning for creating novel view images from single image. Although they produce photorealistic results, they struggle with temporal and spatial consistency, especially in dynamic scenes.
@@ 2.2 Multi-view Novel View Synthesis
Early approaches in this field frequently utilized geometric methods, such as multi-view stereo reconstruction [54, 30] and structure-from-motion techniques [53, 46], to generate novel viewpoints from multiple images captured at different angles. NeRF-based methods were proposed to generate novel views based on fewer input images, including SinNeRF [72], CoPoNeRF [24] and NViST [28] . 3D Gaussian Splatting [32] presented an explicit representation of the scene. Gaussian-based methods, such as pixelSplat [12] and MVSplat [14], can produce high-quality synthesis results from multi-view images. Recently, researchers have begun exploring the application of advanced diffusion models for novel view synthesis, driven by their remarkable success in this area. Notable approaches such as Photoconsistent-NVS [75], MultiDiff [42], GCD [63] and NVS-Solver [74] generate image sequences based on predefined camera trajectories using either single or multiple input images. While these methods have shown promising results, they often struggle with dependence on dense scene geometry, handling complex scene dynamics, and generalization challenges [32]. Moreover, the challenge of optimizing NeRF/GS with sparse views hampers the ability of multi-view NVS to produce high-quality results.
@@ 3 SpatialDreamer
Given an image sequence and stereoscopic pose describing the scene depth information, the proposed method generates a stereo pair. In this section, we first provide a brief introduction about SVD [7]. Secondly, we offer detailed information of our video generation module DVG, which can generate sufficient paired video with geometric and temporal consistency. Using DVG, we present the framework for self-supervised learning-based video synthesis to support efficient and dedicated training. More importantly, to ensure consistency between frames, we design a consistency control module for geometric and temporal consistency ensurance. Finally, we present the detailed training process.
@@ 3.1 Preliminaries
The challenge in video generation lies in accurately modeling the spatial-temporal dynamics, involving spatial relationships within frames and temporal relationships across frames. To address this challenge, Stable Video Diffusion (SVD) [7] introduces temporal convolution and attention layers. SVD follows the latent diffusion model (LDM) to encode video pixels into the latent space, enabling a more efficient denoising process. Given a video $I_{v}$, encoder $\mathcal{E}$ encodes $I_{v}$ into the latent space as $z_{0}=\mathcal{E}(I_{v})$. Gaussian noise $\epsilon$ is then added in the Markov process:
@@
where $\bar{\alpha}_{0},\ldots,\bar{\alpha}_{T}$ are the pre-defined noise schedules for T steps in the Markov process. In particular, SVD is trained by the $v$-prediction formulation with the mean squared error (MSE) between the ground truth and its prediction in the latent space, known as $l_{\epsilon}$.
@@ 3.2 Depth based Video Generation
To create training pairs of images along with their corresponding occluded masks, the following self-supervised mechanism is employed utilizing real images. Firstly, the depth of the given video sample is estimated [49]. Secondly, the reference view image $x_{1}$ is rendered into a masked one under the target viewpoint $P_{2}$, and then the trained inpainting model [58] is used to fill these occluded regions to obtain the novel viewpoint $x_{2}$ [43]. Finally, $x_{2}$ is rendered to the original viewpoint $P_{1}$ via a forward rendering. After this, we can obtain the masked image $\widetilde{x}_{1}$ with mask $M$ under viewpoint $P_{1}$, as well as the inpainted image $x_{2}$ under viewpoint $P_{2}$. The mask $M$ and masked image $\widetilde{x}_{1}$ are considered to be the conditions, with the image $x_{1}$ serving as the ground truth for the diffusion model. This approach involves encoding the inpainted image $x_{2}$ into latent feature representations and feeding it into RefinerNet.
@@ 3.3 Self-Supervised Stereo Video Synthesis framework
Leveraging the generated paired view videos, we propose a self-supervised framework for stereo video synthesis, as illustrated in Figure 2. This framework commences with the rendering of the target view video that incorporates occlusions, followed by the application of a spatio-temporal method to effectively inpaint the missing regions, ultimately yielding a complete non-occluded target view video.
To learn the distribution of the feature space between the paired view images, we devise RefinerNet, which is modeled after the architecture of the denoising U-Net, excluding the temporal layer. RefinerNet adopts the weight initialization from the original SD2.1 model [50]. As depicted in Figure 2, a spatial attention layer is employed instead of the self-attention layer. Given a feature map $z_{t}\in\mathbb{R}^{t\times h\times w\times c}$ from the denoising U-Net and another feature map $z_{r}\in\mathbb{R}^{t\times h\times w\times c}$ from RefinerNet, they are concatenated along the h dimension, and then self-attention is applied. Subsequently, the first half of the feature map is retrieved as the output. One great advantage of this approach is that it enables RefinerNet to maintain the feature representation capability inherent in the original SD Unet, extracting reference view image information. In addition, the denoising U-Net can adaptively learn features from RefinerNet that are correlated in the same feature space, owing to the fundamentally similar network architectures between the denoising U-Net and RefinerNet. This not only ensures a robust initialization for the feature representation but also enhances the learning ability of the entire network.
@@ 3.4.1 Temporal Interaction Learning module
While our RefinerNet-based approach providing spatial guidance, paired view features are still merged independently, posing challenges for temporal consistency and potentially causing texture discrepancies or odd results between adjacent frames.
Inspired by previous studies [11, 22], we propose an attention-based temporal interaction learning module that learns the distribution of the feature space from adjacent frames, named TIL. Specifically, given latent features of $N_{r}$ adjacent images under reference view $z_{i}^{t}$, where $i$ = 1, 2, …, $N_{r}$ and the latent feature of the reference view image $z_{r}^{t}$ at the time step $t$, we augment the $z_{r}^{t}$ to $aug_{r}^{t}$ as:
@@
where $\lambda\in[0,1]$ and $\mathrm{Attn}_{i,j}$ defines attention between two latent features $z_{i}$ and $z_{j}$. This module blends self-attention of $z_{r}^{t}$ in RefinerNet’s U-Net block with the cross-view attention between $z_{r}^{t}$ and each adjacent view $z_{i}^{t}$. The cross-view attention aligns the feature of all adjacent frames to the reference view while the self-attention helps each reference image retain distinctiveness. The augmented reference feature $aug_{r}^{t}$ is then fed into spatial attention layer to assist U-Net network learning.
@@ 3.5 Training Strategy
The training process is divided into two stages. In the first training stage, we focus on individual frames from videos. The temporal layer in denoising U-Net is frozen, and cross-view attention in TIL module is removed. The RefinerNet model and the denoising U-Net are initialized with pretrained weights from SD2.1 [50] and SVD [7] respectively. The weights of the variational autoencoder encoder and decoder, as well as the contrastive language-image pre-training (CLIP) image encoder, are all kept fixed. The aim of this stage is for the model to learn to synthesize new viewpoint images under the condition of a reference image and a new viewpoint pose. In the second stage, we train the temporal layer and cross-view attention in TIL module with video sequences. This enables the model to capture temporal context information efficiently. The input for the model is the 8-frame video clip and $N_{r}$ for TIL is set to 8.
@@ 4.1 Datasets, Baselines and Metrics
We conduct our experiments on image and video levels with RealEstate10K [78] and self-collected video data. RealEstate10K is a large, widely used dataset with camera poses, corresponding to 10 million frames derived from about 80,000 video clips, gathered from about 10,000 YouTube videos. As there is not open-source stereo video benchmark, we collect 1500 monocular videos with 1400 for training, 100 for testing. For image synthesis, we compared the proposed method with recent open-source SOTA methods: 3D-photography [56], SynSin [69], AdaMPI [21], SinMPI [47], Wang et al. [66], Photoconsistent-NVS(P-NVS) [75], NVSVDE-Net [4], CoPoNeRF [24] and MVSplat [14]. For a fair comparison, the 3D photography method, AdaMPI, SinMPI, and the proposed method all used MiDaS [49] for depth estimation. Following the method of 3D-photography [56], we measured the accuracy of the generated target views with regard to the ground-truth images using the three metrics of the learned perceptual image patch similarity (LPIPS), peak signal-to-noise ratio (PSNR), and structural similarity index measure (SSIM).
For the video synthesis, we picked several representative approaches, i.e., the method of 3D-photography [56], the open-source code repository webui-depthmaphttps://github.com/thygate/stable-diffusion-webui-depthmap-script , P-NVS [75], CoPoNeRF [24], NVS-Solver [74] and AVP [44]. Moreover, we conducted the same depth estimation method [49] for a fair comparison of 3D-photography [56], webui-depthmap and NVS-Solver [74]. We employ the FVD [62] score to measure the perceptual similarity between input videos and outputs and report the flow warping error $E_{warp}$ [35] to assess the temporal consistency of the resulting video sequences.
@@ 4.2 Implementation
We employed the AdamW [39] optimizer with a learning rate of $1\times{10}^{-5}$ to train the model on two NVIDIA A800 GPUs. In the first training stage, individual video frames were resized and center-cropped to a resolution of 1024 × 1024. We trained the model for 100,000 iterations with a batch size of 8. In the second training stage, the videos were first split according to the transitions to ensure that a scene only appeared in one video clip. We then formed each video clip for training with eight frames. Finally, the temporal layer and TIL module was trained for 10,000 steps with a batch size of 2. The stereo-aware loss scaling factor $\lambda$ was set to 0.001 and $\lambda$ in Eq. 3 was set as 0.6.